OpenStudy (anonymous):

whats y=2(x+1)(x+4) In standard form

jigglypuff314 (jigglypuff314):

standard form of parabola: y = ax^2 + bx + c so just foil and factor

jigglypuff314 (jigglypuff314):

y = 2(x + 1)(x + 4) = 2( x*x + 4x + 1x + 1*4) = 2(x^2 + 5x + 4) then distribute the 2

OpenStudy (anonymous):

thanks your really helping here , im trying to study for my semester exam , i have algebra as an elective, and lately ive been getting caught up in it pretty bad....